OpenStudy (cutebunny12):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b minus one equals thirty Step 1: fifteen times b minus one minus one equals thirty minus one Step 2: fifteen times b equals twenty-nine Step 3: fifteen times b divided by fifteen equals twenty-nine divided by fifteen Step 4: b equals twenty-nine fifteenths

OpenStudy (lynfran):

15b-1=30 15b=30+1 b=31/15 compare to see where he when wrong...